A beam of light is directed upward through a pool of water (n = 1.33) into air at an angle of 10 degrees from the normal to the water’s surface. At what angle from the normal does the light beam emerge into the air?

a.

10 degrees

b.

7.5 degrees

c.

5.0 degrees

d.

13 degrees

e.

16 degrees
Answer #1

In this case we need to use snell's law:

n_1\sin\theta_1=n_2\sin\theta_2

where 1 is the initial surface (water n=1.33)

1.33\sin(10)=\sin\theta_2

\theta_2=13.35^{\circ}

option d 13 degrees